OpenStudy (anonymous):

Sin4x/sinx=4cosxcox2x

OpenStudy (anonymous):

Use the fact that sin(A+B)=sinAcosB + cosAsinB So sin(2A)=sin(A+A) = 2sinAcosA So sin(4x) = sin(2*(2x)) = 2sin(2x)cos(2x) = 2*(2sinxcosx)cos(2x) = 4 sinx cosx cos(2x)

OpenStudy (anonymous):

Sin4x/sinx 2sin(2x)cos(2x)/sin(x) 4sin(x)cos(x)cos(2x)/sin(x) 4cos(x)cos(2x)